OpenStudy (anonymous):

Automobiles generally depreciate 14% per year. Hamid purchased a Smart Car for $25 000. (5 marks) Create an equation that represents the value of this new vehicle over time, in years. What would its value be after 20 years? How long will it take for the Smart Car to lose three-quarters of its original value?

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Give it a go. Show at least the first year.

OpenStudy (anonymous):

I have no idea where to start :( @tkhunny

OpenStudy (tkhunny):

I never believe that. It simply cannot be true. Have you EVER heard the word "Depreciate"? What does Depreciation do?

OpenStudy (anonymous):

it goes down

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Perfect. If it depreciates 14%, HOW MUCH does it go down?

OpenStudy (anonymous):

by 14%

OpenStudy (anonymous):

0.14

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

It's not necessary to tag me every time. 14% = 0.14 is the RATE that it goes down. What is the AMOUNT that it goes down?

OpenStudy (anonymous):

the smart car and btw I tag you so you see this quicker

OpenStudy (tkhunny):

I see it just as quickly when you post ANYTHING. If you tag, that just means i see it twice. You didn't answer my question. What AMOUNT does the value decrease in the first year?

OpenStudy (anonymous):

25,000

OpenStudy (tkhunny):

No, $25,000 is the total value a the beginning of the 1st year. The value then decreases by 14%. What is 14% of $25,000?

OpenStudy (anonymous):

its' 3500

OpenStudy (tkhunny):

One more time. What AMOUNT does the value decrease in the 1st year?

OpenStudy (anonymous):

I dont get it

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

I dont understand what you mean

OpenStudy (tkhunny):

$25,000 Depreciates 14% annually. $3500 is 14% of $25,000 What is the value after $25,000 depreciates $3,500?

OpenStudy (anonymous):

ohhh 21500

OpenStudy (anonymous):

what is the equation

OpenStudy (anonymous):

??

OpenStudy (tkhunny):

25000 - 3500 = 21500 NOw, use the EXACT same PATTERN to calculate the AMOUNT of depreciation in the 2nd year and the total value after the 2nd year.

OpenStudy (anonymous):

I did it three times

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

HOW DO I FORM AN EQUATION FROM HERE ??????

OpenStudy (tkhunny):

You must answer my question. NOw, use the EXACT same PATTERN to calculate the AMOUNT of depreciation in the 2nd year and the total value after the 2nd year. It doesn't do any good to tag me if you are not going to answer my questions. We need to get a feel for how it works. This is how we LEARN to formulate equations.

OpenStudy (anonymous):

THE VALUE WOULD BE $18490 AFTER THE SECOND YEAR

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Okay, now one more thing before building the equation becomes natural. EXACTLY how did you obtain the correct answer $18,490 from $21,500?

OpenStudy (anonymous):

I multiplied 21500 by 0.14

OpenStudy (anonymous):

Now can we make the equation? @tkhunny

OpenStudy (tkhunny):

That's not correct. 21500*0.14 = 3010 I don't see 18490, yet.

OpenStudy (anonymous):

oh then I subtracted 21500-3010

OpenStudy (tkhunny):

Okay, here we go. We have: Year 1 25000 * 0.14 = 3500 25000 - 3500 = 21500 Year 2 21500 * 0.14 = 3010 21500 - 3010 = 18490 Do you agree that this is the complete solution for the first two years?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Can you rewrite both of those years in such a way that we do NOT see the exact amount of the depreciation? Somehow, combine the two steps of the process into a single step that goes straight to the final amount?

OpenStudy (anonymous):

idk how to

OpenStudy (anonymous):

come on help me

OpenStudy (tkhunny):

I'll give you five minutes to think about it and produce a possible result.

OpenStudy (anonymous):

25000x - 0.14

OpenStudy (anonymous):

idk what Im doing

OpenStudy (tkhunny):

That's pretty good. Not quite, but you are in the neighborhood. Year 1 25000 * 0.14 = 3500 25000 - 3500 = 21500 Look at that bottom equation. 25000 - 3500 = 21500 Can you think of another expression for 3500?

OpenStudy (anonymous):

3500x -0.14 ??

OpenStudy (anonymous):

its wrong I know

OpenStudy (tkhunny):

That makes no sense. Look at the first equation. 25000 * 0.14 = 3500 Can you see TWO ways to write 3500?

OpenStudy (anonymous):

ya

OpenStudy (tkhunny):

Okay, now rewrite this equation using some other expression for 3500. 25000 - 3500 = 21500

OpenStudy (anonymous):

so wait is that the equation you wanted me to write ?

OpenStudy (tkhunny):

Rewrite this equation using some other expression for 3500. 25000 - 3500 = 21500

OpenStudy (anonymous):

2500- 1000=21500

OpenStudy (tkhunny):

What is that? 2500 - 1000 = 1500 - That's not 3500 21500 is not 3500. Rewrite THIS equation using some other expression for 3500. Keep in mind that we know that 3500 = 25000 * 0.14 25000 - 3500 = 21500

OpenStudy (tkhunny):

I see. It's a typo. Somehow, you meant 25000 - 1000 = 21500. This, of course, makes no sense, since 25000 - 1000 = 24000. Any way... 25000 - 3500 = 21500 From the other equation, we have 25000 - (25000*0.14) = 21500 This needs to make sense.